Parth (parthkohli):

Quick tutorial on difference of squares. Let's start by multiplying \(a + b\) and \(a - b\) \( \color{Black}{\Rightarrow (a + b)(a - b) }\) \( \color{Black}{\Rightarrow a(a - b) + b(a - b) }\) \( \color{Black}{\Rightarrow a^2 - ab + ab - b^2 }\) \( \color{Black}{\Rightarrow a^2 - b^2 }\) Do you see that? well this is just a difference of squares. Similar for \(p^2 - q^2\) and \(x^2 - y^2\) etc. etc. We can conclude that \((a + b)(a - b) = a^2 - b^2\) More questions accepted below.

Parth (parthkohli):

Questions?

Parth (parthkohli):

If we have a quadratic, we may use the zero product rule.

Parth (parthkohli):

If you guys need more examples: \( \color{Black}{\Rightarrow 9x^2 - 4 }\) \( \color{Black}{\Rightarrow 9x^2 - 2^2 }\) \( \color{Black}{\Rightarrow (3x)^2 - 2^2 }\) \( \color{Black}{\Rightarrow (3x + 2)(3x - 2) }\)

OpenStudy (anonymous):

Love me some squares

Parth (parthkohli):

What if we have \(16x^2 - 36 = 0 \)? We conclude this: \((4x - 6)(4x + 6) = 0\) and then use the product rule. :)

Parth (parthkohli):

I'll do another small tutorial on the product rule soon anyway.

myininaya (myininaya):

Can you write the as a difference of squares: \[-x^4-3x^2+2x-3\] hehe :)

Parth (parthkohli):

Factor by grouping I think ^

myininaya (myininaya):

That's right :)

Parth (parthkohli):

Is that a 2 or a 3?

myininaya (myininaya):

You should be able to write it like this: \[x^2+2x+1-(x^4+4x^2+4)\]

Parth (parthkohli):

And now factor :)

myininaya (myininaya):

I cheated though I worked backwards just to make something hard for you

Parth (parthkohli):

\( \color{Black}{\Rightarrow x^2 + 2x + 1 - (x^4 + 2x^2 + 2x^2 + 4) }\) \( \color{Black}{\Rightarrow x^2 + 2x + 1 - (x^2(x^2 + 2) + 2(x^2 + 2)) }\)

Parth (parthkohli):

Ah

Parth (parthkohli):

\( \color{Black}{\Rightarrow (x + 1)^2 - (x^2 + 2)^2 }\)

myininaya (myininaya):

Yep yep :)

Parth (parthkohli):

And then apply that...

OpenStudy (lgbasallote):

how about writing \[36x^2 - 252\] as a difference of two squares? i cant get how to do it :/

OpenStudy (anonymous):

\[36x ^{2}-36*7=(6x )^{2}-(6\sqrt{7})^{2}\] \[(6x-6\sqrt{7})(6x+6\sqrt{7})\] i think

OpenStudy (theviper):

Good tutorial @ParthKohli

OpenStudy (anonymous):

how about factoring \(x^4+4\)?

OpenStudy (theviper):

\[\color{red}{x^4+4+4x^2-4x^2}=\color{blue}{(x^2+2)^2-4x^2.}\]\[=\color{purple}{(x^2+2)^2-(2x)^2.}\] Now it is the\[\color{green}{a^2-b^2.}\]form, as u r seeing it:) now factorize^_^

OpenStudy (theviper):

I know satellite already know this but it is for parthkohli; if & only if he don't this thing:)

Parth (parthkohli):

lol how does sat dominate my tutorials :P

Parth (parthkohli):

@TheViper woah sat gave you a medal

OpenStudy (campbell_st):

what about \[\sqrt[3]{x^{10}} - \sqrt[5]{y^6}\]

OpenStudy (anonymous):

You just take roots again. ;_;

Parth (parthkohli):

lol I don't know this advanced stuff. Though I know how to express these in exponents. I doubt that will help.

Parth (parthkohli):

They ain't perfect squares though :/

OpenStudy (campbell_st):

ummm I wonder....

OpenStudy (campbell_st):

its not that advanced...

OpenStudy (anonymous):

It isn't. It's just not pretty. :P

OpenStudy (campbell_st):

\[(\sqrt[3]{x^5}-\sqrt[5]{y^3})(\sqrt[3]{x^5} + \sqrt[5]{y^3})\]

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large x^{10 \over 3} - y^{6 \over 5} }\)

OpenStudy (campbell_st):

\[\lim_{x \rightarrow 9} \frac{ 3 - \sqrt{x}}{x - 9}\] was a question asked yesterday

Parth (parthkohli):

Ahhh NO I DON'T KNOW LIMITS!

Parth (parthkohli):

Maybe I do yes... We'll try it with 8.9 then 8.99 then 8.999

OpenStudy (campbell_st):

but you know how to factorise or expand... you don't need to now limits

OpenStudy (anonymous):

He's saying you can use this tutorial on that problem.

OpenStudy (campbell_st):

thanks limitless

Parth (parthkohli):

Yeah. It'd look something like \(\Large x^3 \times x^{1 \over 3} - {y^{1 \over 5} \times y}\)

OpenStudy (campbell_st):

no... I posted the answer above...

OpenStudy (anonymous):

You're welcome, Campbell.

Parth (parthkohli):

Oh yep. Got it :D

Parth (parthkohli):

lol how couldn't I do that *facepalm*

OpenStudy (campbell_st):

my point is parth... there are lots of equations reducible to quadratics.... provided you know what you are looking at

Parth (parthkohli):

Hmmm

OpenStudy (campbell_st):

and the limit question is another example...

Parth (parthkohli):

Shall I close the tutorial now?

OpenStudy (campbell_st):

lol... its you tutorial.... I'm just making comment

OpenStudy (campbell_st):

and the limit quest\[\lim_{x \rightarrow 9}\frac{-(\sqrt{x} -3)}{(\sqrt{x} -3)(\sqrt{x}+3)}..===> \lim_{x \rightarrow 9}\frac{-1}{\sqrt{x}+3}\]ion